OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (6,2) and a directrix of y=1.

OpenStudy (anonymous):

I know how to do these, but this question is throwing me off because the halfway point is 1/2.

OpenStudy (amistre64):

what is the x value of the vertex? the y value is 1/2 as you say

OpenStudy (anonymous):

@amistre64 you mean the 6?

OpenStudy (amistre64):

the 6 is good :) so we have a vertex at 6,.5 this is where i tend to have to do this a different way. i know that the line y-.5 = 1/2 (x -6) has to equate to y = m(x-6)^2-.5

OpenStudy (amistre64):

i just tend to forget the geometric parts of it is all

OpenStudy (anonymous):

Okay well I called my teacher (she finally answered) and she showed me how to do it. I have another one that I know how to work out and it's a little easier. But I might need help for the second part.

OpenStudy (amistre64):